### v1

# Convergence of body-orders in linear atomic cluster expansions

Abstract
We study the convergence of a linear atomic cluster expansion (ACE) potential [1] with respect to its basis functions, in terms of the effective two-body interactions of elemental Carbon systems. We build ACE potentials with descriptor sets truncated at body-orders $K=2,3,4$ trained on pure dimers, or on large datasets of different diversities but without any dimers. Potentials trained on a more diverse dataset fare better in validation and result in a nontrivial dimer curve, but still very far from the theoretical two-body interaction calculated by DFT. Moreover, dimer curves between descriptor sets clipped at different $K$ do not seem to converge to a universal function for a given dataset. We conclude that machine learning potentials employing atomic cluster expansions optimize losses at low $K$ but fail to generalize and converge properties described by two-body interactions.
I Introduction
Decompositions of the total energy of atomic systems by $N$-body contributions, or cluster expansions,
$E\left(\{\mathbf{r}_{i}\}\right)$ $=V^{(0)}+\sum_{i}V^{(1)}\left(\mathbf{r}_{i}\right)+\sum_{\begin{subarray}{c}i,j\\ i\neq j\end{subarray}}V^{(2)}\left(\mathbf{r}_{i},\mathbf{r}_{j}\right)+$
$\sum_{\begin{subarray}{c}i,j,k\\ i\neq j\neq k\neq i\end{subarray}}V^{(3)}\left(\mathbf{r}_{i},\mathbf{r}_{j},\mathbf{r}_{k}\right)+...$ (1)
have been used as the framework to study material properties such as the thermodynamics of substitutional alloys [2, 3] and empirical potentials of condensed matter and molecular systems [4, 5, 6]. While the infinite expansion has been shown to be complete, its practical effectiveness is limited to cases where the total energy converges after a few body-orders $\nu$, as the number of terms increases combinatorially as higher orders are included.
More recently, the expansion has been used in describing local chemical environments for machine learning potentials. The underlying assumption for these descriptors is that the energy of the system is a sum of its individual atomic contributions $E_{i}$ [7],
$E\left(\{\mathbf{r}_{i}\}\right)$ $=\sum_{i=1}^{N}E_{i}\left(\{\mathbf{r}_{i}\}\right)$ (2)
where $N$ indicates the total number of atoms in the structure. The atomic energies may then be expressed as an expansion in the number of neighbors.
Developments on improving the construction of $E_{i}$ have been made in recent decades, with symmetry-respecting descriptors that use only a few terms in the expansion [7, 8, 9, 10] or take advantage of tensorial operations to reach an arbitrary body order [11, 1, 12, 13]. A detailed review on descriptors of local chemical environments has been done by Musil et al. [14]
While the completeness [2, 3, 15] of such expansions allow an apt description of the system, its form in Eq. 1 are considered to converge slowly [^1] Counterarguments [16] against the speed of convergence of body-ordered expansions claim that the surrounding environments are typically ignored and are usually crucial in the convergence at low body-orders. [4, 5, 17], where computing a term of order $\nu$ would scale in powers of $\nu-1$ of the number of neighbors. In order to avoid this prohibitive scaling, it is possible to express higher body-order terms as the product of simple two-body terms, at the price of including lower order terms, and losing the orthogonality of the expansion. A popular version of this resummation is the Atomic Cluster Expansion proposed by Drautz [1]. In this approach, each $E_{i}$ is expanded with a non-orthogonal basis, $\{A_{\nu}\}$, with unrestricted sums,
$E_{i}\left(\{\mathbf{r}_{i}\}\right)$ $=\sum_{\nu}c^{(1)}_{\nu}A_{i,\nu}+\sum_{\nu_{1}\nu_{2}}^{\nu_{1}\geq\nu_{2}}c_{\nu_{1}\nu_{2}}^{(2)}A_{i,\nu_{1}}A_{i,\nu_{2}}+...$ (3)
with single sums over neighbors
$A_{i,\nu}=\sum_{j}f_{\nu}\left(\mathbf{r}_{i},\mathbf{r}_{j}\right),$ (4)
and $\nu$ indexing the terms of the expansion.
The expansion has been shown to be complete [15], and implementations of the code have shown promise in machine learning applications [18, 19, 20, 21]. One of the state-of-the-art potentials, MACE [13], extends from the formalism of ACE and represents the state of the art in molecular dynamics applications for materials with multiple elements [22].
Whereas potentials adopting ACE descriptors have shown to converge fast in losses, a detailed study on the convergence of the body-orders with respect to properties described by lower orders in the original expansion in Eq. 1 have not been thoroughly investigated. In this work, we investigate the convergence of linear ACE at different body-orders $K$ in terms of their projection on the lowest terms of the original cluster expansion Eq. 1. In particular, we consider the convergence of the potential in energy and forces on different datasets, as a function of the number and type of basis functions utilized, and compare it to the stability of a real body ordered expansion. Our study focuses on the first nontrivial term of the expansion: the two-body term. This can be easily extracted from a given potential by studying the energy profile of two isolated atoms, a dimer, as a function of distance:
$E_{2b,i}\left(\mathbf{r}_{i},\mathbf{r}_{j}\right)$ $=V^{(2)}\left(\mathbf{r}_{i}-\mathbf{r}_{j}\right).$ (5)
It is thus natural to ask whether this function might converge to the real dimer dissociation curve, or to another arbitrary function for a dataset that does not contain dimers.
Neural network potentials (NNPs) may also employ nonlinearities via functional embeddings [23, 24, 6], or are inherent from their activation layers[7, 25, 9, 10, 14]. While nonlinearities improve their performance by being able to represent higher-order terms when the expansion is truncated, one has to be careful in how they are applied as certain applications may invalidate the body-ordering of the descriptor [12]. We do not aim to build the most accurate potentials, but to have a more grounded understanding of their behavior in the ACE framework. As such, we limit our investigations to training simpler linear models as well as not going to too high of a body-order where returns are diminished. Linear ACE models perform comparably well relative to neural network potentials, and have been successfully employed to describe molecular [18] and condensed phases [19].
II Methods
II.1 Datasets
The datasets we use in the investigations consists of elemental Carbon structures. Three datasets were used: the first set consists of $247$ dimer structures at logarithmically-spaced distance between $0.5$ and $8$ Angstrom. Half of the structures were randomly selected for the validation set. The remaining two are sets of $1000$ structures coming from a collection of Carbon allotropes generated from [26]. One subset contains a homogeneous collection of diamond-like structures clustered using a Euclidean distance measure [27]. The diamond-like dataset is highlighted in the histogram of distances to a reference diamond structure in Fig. 1. The last subset are structures obtained using farthest point sampling [28] to maximize its diversity. For the diamond-like and diverse datasets, a validation set was constructed from $1000$ diverse structures that have not been seen in training. All structures studied energies and forces were calculated using the Quantum Espresso package [29] using the rVV10 functional [30] to account for nonlocal interactions.
Figure 1: Histogram of fingerprint distances of the configurations in the Carbon dataset in [26] relative to a reference diamond-like structure (inset). $1000$ configurations were chosen within a narrow region in structure space.
II.2 Descriptor
We use the ACE descriptor as implemented by the pacemaker package [20]. We used scaled Chebyshev polynomials as the radial functions, vanishing smoothly at the cutoff length of $r_{\text{cut}}=5$. We start from $K=2$ body functions and progressively add more terms of the same order. We then introduce higher order $K=3$ terms while keeping the number of lower order terms fixed, and then repeat the procedure for higher orders. ACE descriptors employ spherical harmonics to describe angular functions, with $n,l$ values chosen to respect maintain rotational symmetry. Low $n,l$ values were chosen to avoid overfitting the model from too many basis functions. We note that the original ACE and the current version of pacemaker include self-interactions from $K>2$ body orders, so redundancies of lower-body order functions appear as functions of higher $K$.
II.3 Fitting procedure
We use the default loss function $\mathcal{L}$ of pacemaker [20],
$\mathcal{L}=\left(1-\kappa\right)\Delta^{2}_{E}+\kappa\Delta^{2}_{F}+\Delta_{\text{coeff}}+\Delta_{\text{rad}}$ (6)
where $\Delta^{2}_{E}$ and $\Delta^{2}_{F}$ are the corresponding energy and force root-mean-square errors (RMSE), and $\kappa$ is the coefficient that balances the energy and force contributions to $\mathcal{L}$. The term $\Delta_{\text{coeff}}$ regularizes the basis function coefficients [^2] This term includes the L1 and L2 regularizations., while $\Delta_{\text{rad}}$ regularizes the coefficients of the radial components of the basis function. The dimer dataset was randomly split between training and validation set, while the diamond-like and generic datasets were validated using a set similar to the generic set. To reconstruct the 2-body contribution to the cluster expansion, we build the energy curve of two isolated Carbon atoms as in Eq. 5. The converged potentials from pacemaker were evaluated using the Atomic Structure Environment (ASE) code [31].
III Results and Discussion
Figure 2: Top: $\mathcal{L}$ as a function of number of basis functions, $N_{f}$, for the dimer dataset. Training and validation errors as dashed and solid curves. Curve colors indicate potentials built from basis sets which are exclusively two-body (light green), up to three-body (green), and up to four-body functions (dark green), respectively. Light solid lines indicate $\mathcal{L}$ if more functions introduced keeping the previous $K=2$ or $K=3$. Bottom: dimer curves for each of the potentials trained at certain $N_{f}$, with reference DFT data as points. Heavy lines represent the potential with the largest $N_{f}$ before functions of higher $K$ are introduced.
We start by fitting the dimer dataset with an increasing number of basis functions, $N_{f}$, up to $250$ elements and $K=4$ body-order. The loss and dimer curves are shown in Fig. 2, where light green, green, and dark green indicate data from potentials trained using $K$ up to 2, 3, and 4, respectively. The top plot shows the dependence of the loss $\mathcal{L}$ from Eq. 6 to $N_{f}$. The validation loss (solid curve) converges $K=2$ and increases with the addition of higher order functions, jumping two orders of magnitude as more $K=4$ functions are added. This is of course to be expected, as the dataset consists solely of 2 body structures. The bottom plot highlights the dimer curves of the potentials trained from the dimer set at the highest $N_{f}$ of a particular order $K$. We see that all potentials were able to reproduce the dimer curve as expected from the DFT results.
The energy RMSE for the representative potentials were around $5$ meV/at, and force RMSE were slightly less than $200$ meV/Å for $K=2$ and $K=3$. The force RMSEs increased to $2700$ meV/Å for the overfitting potential that at $N_{f}=250$, and are due to the short-range region of the curve having energies up to almost two orders of magnitude higher, making the deviations from DFT results more sensitive to errors. A comparison of the forces at the short range region may be found in the Supporting Information B.1.
The results for the homogeneous diamond-like dataset are shown in Fig. 3. Unlike the potentials trained from the dimer dataset, the training loss from the diamond-trained potentials reduce by an order of magnitude upon the introduction of the $K=3$ functions. The light orange line indicates that given the same $N_{f}$, the basis set purely constructed from $K=2$ functions does not have as pronounced of a decrease in $\mathcal{L}$. We see that the validation over the diverse set has losses improving for $K=3$, but overfitting occurs at $K=4$. The dimer curves from these potentials show a considerable amount of variability even after the loss is converged, and they remain far from the expected dimer behavior dictated by DFT. Of course, the lack of diversity in this dataset, leading to very small variations in the radial distribution function, might be part of the reason why a rather flat 2-body curve is obtained. Moreover, the overfitting for the large descriptor size could contribute to the variability of this contribution.
Figure 3: $\mathcal{L}$ vs. $N_{f}$ (top) and dimer curves (bottom) for the homogeneous, diamond-like dataset. Orange and pink plots correspond to potentials with only $K=2$ and up to $K=3$ functions, respectively while purple curves include $K=4$ functions. The x-values of the dimer curve start at the minimum distance found in the homogeneous dataset.
We then look at the dimer curves produced by the diverse dataset for different descriptor complexities, ranging from $K=2$ to $5$, with various sizes for $K=4$ (similar loss analysis with consistent earlier dimer curves are shown in the Supporting Information B.2.). Fig. 4 shows these curves having more pronounced shapes as compared to the ones produced by the diamond-like dataset. While the curves look similar as more functions of the same $K$ are introduced in the set as seen in the $K=4$ potentials, the curves of different $K$ limits do not seem to converge even when five-body contributions are already added. The variability is much more pronounced, even among potentials with similar maximum $K$ and more so when adding new terms. We posit that the effective two-body potential does not seem to converge, even for when five-body basis functions are added, leading to an arguably well-converged potential.
, Refer to caption
Figure 4: Dimer curves for the diverse $1000$-structure dataset for extended descriptor sizes. While curves produced by models within a certain $K$ ($K=4$) may appear to converge, they qualitatively do not seem to approach a consistent shape whenever functions of higher $K$ are introduced in the descriptor set.
Lastly, we train a potential that includes the $50000$ structures on the whole Carbon dataset in [26] to see if further increasing the complexity of the input data affects our conclusion. The dimer curves shown in Fig. 5 tell the same story as the one of smaller dataset: the curves do not approach the DFT dimer results at all, and they do not seem to converge to a fixed function as higher $K$ basis elements are introduced.
It is worth reminding that we are testing the generalization capabilites of the diverse set potentials and not their ability to capture in-domain data. The pair-interactions present in the set are those that exist in surrounding materials and not in isolation. A brief test shown in Supporting Information B.3. shows that gradually populating the diverse dataset with dimers eventually leads to a binding curve that approaches the DFT results.
Figure 5: Dimer curves for expanded $50000$-structure dataset for multiple descriptor sizes. Like the potentials trained with a smaller dataset in Fig. 4, potentials trained with increasing descriptor sizes but within a certain $K$ seem to converge, but introducing the descriptor set with $K=5$ basis functions vary the shape considerably.
Table 1: Comparison of linear ACE models to neural network potentials in MAE for energies and forces for the Carbon dataset in [26] for datasets containing $n_{\text{set}}=1000$ and $50000$ structures. Energies (E) are in meV/atom and force components (F) are in meV/Å.
$n_{\text{set}}=1000$ $n_{\text{set}}=50000$
Potentials E F E F
$K=4,250N_{f}$ 48.50 241.27 44.06 222.42
$K=4,493N_{f}$ 18.30 199.23 17.55 176.48
$K=4,930N_{f}$ 15.40 167.73 15.27 166.70
$K=5,793N_{f}$ 14.88 160.78 14.74 159.80
PANNA small 22.40 248.85 17.76 162.37
PANNA mid 20.93 241.36 12.41 137.45
PANNA big 17.88 242.02 8.21 166.95
NequIP $l=1$ 20.40 213.00 7.96 110.00
NequIP $l=2$ 6.83 105.00 5.21 51.40
MACE 6.81 103.14 1.82 51.31
We also make sure that the performance of our linear models are reasonable by comparing their loss metrics to the performance of different neural network potentials trained on the same dataset in [32] in Table 1. For the $1000$-size subsets, we found that the linear ACE models with the most complex descriptor sets ($K=4,N_{f}=930$ and $K=5,N_{f}=793$) were able to outperform all PANNA potentials and even the $l=1$ version of the NequIP [33] potential. However, they perform relatively worse with all other potentials when the complete $50000$-size Carbon set is used as the training set, which suggests that the nonlinearities absent in the ACE models provide a considerable improvement in training more complex datasets. Despite this, we assert that the inability of the linear models to capture the effective two-body interaction and converge to a defined shape are not due to the poor $\mathcal{L}$, but rather the nature and truncation of the expansion employed in the descriptor sets.
IV Summary and conclusions
In this work, we studied the convergence of linear cluster expansions with respect to the $N$-body basis functions through the effective two-body interactions, $E_{2b}$, of elemental Carbon systems. We used the atomic cluster expansion (ACE), with descriptor sets truncated at body-orders $K=2,3,4$ trained on pure dimers, then on two datasets of different diversities but without any dimers. The potentials trained on the dataset limited to diamond structures generally performed poorly: overfitting on training data, and producing a flat dimer curve. The potential trained on a more diverse dataset fared better in the validation errors and resulted in a nontrivial curve dimer curve, but still very far from the theoretical two-body interaction calculated by DFT. Moreover, the dimer curves between descriptor sets clipped at different $K$ did not seem to converge to a universal function for a given dataset. Even training on the full original dataset led to the same results. We conclude that machine learning potentials employing linear cluster expansions optimize losses at low $K$, but fail to generalize and converge properties described by two-body interactions.
A recent study in  [34] provides a canonical ACE where self-interactions are removed. Their tests on properties from out-of-domain elements show that while the potentials trained with canonical ACE provides smoother dimers as compared to the ones trained with the self-interacting ACE, the potentials at different regularization tolerances do not stabilize to a single curve. This suggests that the inability of the potentials to converge properties at low $K$ may go beyond the contributions of the self-interactions and is innate in the expansion itself.
V acknowledgement
This study was funded by the European Union – NextGenerationEU – PNRR M4C2-I.1.4, in the framework of the project CN-HPC: National Centre for HPC, Big Data and Quantum Computing - Spoke 7 - (MUR ID: CN_00000013 – CUP: G93C22000600001). FP work was supported by EU – FSE REACT-EU and Italian MUR under contract 45-I-D2100-1 of the PON-RI 2014-2020 initiative. SdG work was supported in part by the European Commission through the MAX Centre of Excellence for supercomputing applications (grant numbers 10109337 and 824143). Computational resources were provided by CINECA.